Hi! I just want to know how to transform this equation to a quadratic equation.. and Kindly please tell me how you did it. Just for my physics problem. Thanks! :D
-40 = 100t + 1/2 (-9.8)t^2
It should come like this, right?
-40 = 100t - 4.9t^2
But how to arrange them? thanks! :D

Answer by rothauserc(4718)

-40 = 100t + 1/2 (-9.8)t^2
-4.9t^2 +100t +40 = 0
this is an equation of a parabola that opens downward
max height in meters is determined by solving for the axis of symmetry
t = -b/2a = -100 / (2*-4.9) = 10.204081633, therefore
max height occurs at 10.204081633 seconds and max height is
h(t) = -4.9(10.204081633^2) +100*(10.204081633) +40
max height = 550.204081633 meters